OpenStudy (anonymous):

The circumference of a circle is 19.468 meters. Find the radius of the circle. 3.1 m 6.2 m 16.328 m 12.4 m

OpenStudy (michele_laino):

the circumference C of a circle is given by the subsequent formula: \[\Large C = 2\pi R \cong 6.28R\] where R is the radius of your circle

OpenStudy (anonymous):

6.2

OpenStudy (anonymous):

B.

OpenStudy (michele_laino):

if we divide both sides by 6.28, we get: \[\Large R = \frac{C}{{6.28}} = \frac{{19.468}}{{6.28}} = ...?\]

OpenStudy (anonymous):

3.1

OpenStudy (michele_laino):

correct!